OpenStudy (anonymous):

a rectangle has an area of 96 sq ft if the width of the yard is 4 ft less than the length, what is the perimeter in ft of the yard: that's my problem.. my question is what is the number that will help me figure this out or how do i get that number to find my answer

OpenStudy (anonymous):

put say \(x\) as the length, then \(x-4\) is the width because it says the width of the yard is 4 ft less than the length

OpenStudy (anonymous):

the area is length times width, and so you have to solve \[x(x-4)=94\] can you do that?

OpenStudy (anonymous):

oh ok yes i can

OpenStudy (anonymous):

ok then once you find \(x\) you still have to find the perimeter, which will be \[x+x+x-4+x-4\] better known as \[4x-8\] don't forget that step at the end

OpenStudy (anonymous):

let me know what you get for \(x\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

44 my final answer

OpenStudy (anonymous):

hmmm i think i got 40 what did you get for \(x\) ?

OpenStudy (anonymous):

dammit well it doesnt matter ur correct

OpenStudy (anonymous):

i used 24

OpenStudy (anonymous):

oh damn i made a mistake above, it is \(96\) not \(94\) you have to solve \[x(x-4)=96\]

OpenStudy (anonymous):

\[x(x-4)=96\] \[x^2-4x=96\] \[x^2-4x-96=0\]\[(x-12)(x+8)=0\]

OpenStudy (anonymous):

then if \(x=12\) you have \(4\times 12-8=40\)